OpenStudy (anonymous):

How do you find the natural frequency of a pendulum that has a length of 1.0m (assuming that gravity is 9.8m/s^2)

OpenStudy (calculusfunctions):

frequency = 1/period The period for a pendulum is given by\[T =2\pi \sqrt{\frac{ L }{ g }}\]where L is the length of the string attached to the pendulum bob, g is the acceleration due to gravity, and T of course is the period of the pendulum. Find the reciprocal of the period, to find the frequency.

OpenStudy (theraggedydoctor):

c;